# MACE-OFF: Short Range Transferable Machine Learning Force Fields for Organic Molecules

### v3

@@ III.6.4 Protein simulation in explicit solvent
The results of sections III.5 and III.6 highlight the most significant differences between previous transferable organic ML potentials and the MACE-OFF models. These are, to our knowledge, the first local ML potentials capable of reproducing experimental properties of condensed phase molecular systems, as well as large-scale, all-atom, explicit solvent biomolecular simulations without relying on augmented, system-specific training data.
@@
Figure 13: Protein in explicit solvent. Power spectrum of the high hydration state of crambin computed from the final 125 ps of a 1.6 ns simulation with MACE-OFF24(M). GEMS, AMBER and SO3LR spectra are reproduced from Ref. [41].
III.7 Computational performance
@@
To evaluate the computational performance of the MACE-OFF models, we performed NVT molecular dynamics for water using both OpenMM and LAMMPS, with simulation boxes ranging from $10^{2}$ to $10^{5}$ atoms. The OpenMM results shown in Figure 14 were collected with a single NVIDIA A100 80GB GPU. The LAMMPS results shown in Figures 15-16 highlight both CPU and GPU evaluation.
@@
In addition to the native PyTorch implementations of the MACE-OFF23 models, we also report results for optimized implementations leveraging custom C++, CUDA, and/or Kokkos code. One optimized variant, MACE-OFF23(S)-OPT, incorporates custom CUDA kernels which are packaged as a standalone, inference-only library, cuda_mace [92]. Another pure-C++ implementation, as well as a Kokkos implementation, will be made available as part of the symmetrix library [93].
@@
In these optimized variants, the learnable radial function $R^{(t)}_{k\eta_{1}l_{1}l_{2}l_{3}}(r_{ij})$ is evaluated with cubic splines, which efficiently map $r_{ij}$ to each of the outputs of the radial MLP. Furthermore, Eqs. II.1 and 3 are implemented in unified kernels that simultaneously compute the one-particle basis and perform the summation over neighbors to generate the invariant atomic basis $A_{i}$. The computation of spherical harmonics, as detailed in equation II.1 is accelerated using the sphericart [94] library, which offers a highly optimized implementation for real spherical harmonics. All linear and equivariant linear operations have been substituted with custom kernels, and the cuda_mace implementation further leverages tensor core acceleration while incorporating error-correction techniques to ensure numerical accuracy. Additionally, the computation of the product basis $\mathbf{B}_{i}$ in equation 4 is optimized using the sparsity of the generalized Clebsch-Gordon coefficients $\mathcal{C}^{LM}_{\eta_{\nu}\bm{l}\bm{m}}$, significantly accelerating the product calculations over both the body order $v$ and the coupling terms.
@@
Figure 14: Performance in OpenMM. Performance of MACE-OFF23 in OpenMM for liquid water in the NVT ensemble at 1 g/cm3 at 300K with a 1 fs timestep on a single NVIDIA 80GB A100 GPU.
Figure 15: Throughput in LAMMPS. Performance of molecular dynamics in LAMMPS for water in the NVT ensemble at 1 g/cm3 and 300 K with a 1 fs timestep. All simulations used variants of the MACE-OFF23(S) model.
The left panel highlights CPU performance achievable on the UK-based Archer2 CPU supercomputer, using the new pure-C++ MACE implementation. The solid black curve demonstrates the throughput of a single 128-core node as a function of the number of atoms. The two dotted curves demonstrate weak scaling with 432 atoms/node or 3600 atoms/node. All calculations used double precision floating point numbers.
The right panel demonstrates the performance of the GPU evaluators. The three solid curves provide single-GPU results (determined with NVIDIA A100 80GB GPUs) for the native Torch-based MACE-OFF23(S), a custom Kokkos-based MACE implementation, and the cuda_mace implementation marked as OPT*. The asterisk indicates that these latter simulations used single precision floating point numbers, with error correction for the most numerically sensitive terms, in contrast to the remaining calculations which were performed with double precision. Finally, the two dashed curves demonstrate the performance when multiple GPUs are employed, using the pure-Kokkos implementation, to access larger systems with domain decomposition. More work is needed to improve this weak scaling.
Figure 16: Evaluation speed in LAMMPS. Data from Figure 15, replotted to show the average duration of each timestep.
@@ IV Outlook
In this paper, we presented the series of MACE-OFF force field and dipole models, demonstrating the broad applicability, transferability and capability of purely short-range machine learning potentials for organic (bio)molecular simulations. We have shown that the models, based on the MACE higher order equivariant message passing architecture, can improve on the pioneering ANI models [27] that served as the only widely applicable machine learning force fields for molecules for a number of years. While a full ablation study between different ML models is not possible due to a wide range of design choices made in their construction, the MACE-OFF model architecture, training data and loss function all contribute to significant improvements in both accuracy and extrapolation, combined with high computational speed.
@@ Acknowledgements
The data used to train the models are publicly available at: https://doi.org/10.17863/CAM.107498. The torsion drive dataset is also available at: https://zenodo.org/records/11385284. The MACE-OFF series of models are available at: https://github.com/ACEsuit/mace-off/tree/main and the MACE-OFF23-$\mu$ model is available at: https://github.com/venkatkapil24/MACE-OFF-mu-models

### v4

@@ III.4.4 Protein simulation in explicit solvent
The results of sections III.3 and III.4 highlight the most significant differences between previous transferable organic ML potentials and the MACE-OFF models. Whilst there is still significant work to be done to comprehensively show improvement over classical forcefields for large biomolecular systems, these results demonstrate the capabilities of models fit ‘bottom up’ from quantum mechanical data to learn subtle intermolecular interactions and correctly predict macroscopic properties.
@@
Figure 11: Protein in explicit solvent. Power spectrum of the high hydration state of crambin computed from the final 125 ps of a 1.6 ns simulation with MACE-OFF24(M). GEMS, AMBER and SO3LR spectra are reproduced from Ref. [41]. Black lines denote experimental peaks at 1640 cm-1 and the range 3200-3600 cm-1.
III.5 Computational performance
@@
To evaluate the computational performance of the MACE-OFF models, we performed NVT molecular dynamics for water using both OpenMM and LAMMPS, with simulation boxes ranging from $10^{2}$ to $10^{5}$ atoms. The OpenMM results shown in Figure 12 were collected with a single NVIDIA A100 80GB GPU. The LAMMPS results shown in Section S9 highlight both CPU and GPU evaluation.
@@
In addition to the native PyTorch implementations of the MACE-OFF23 models, we also report results for optimized implementations leveraging custom C++, CUDA, and/or Kokkos code. One optimized variant incorporates custom CUDA kernels which are packaged as a standalone, inference-only library, cuda_mace [88]. Another pure-C++ implementation, as well as a Kokkos implementation, will be made available as part of the symmetrix library [89].
@@
In these optimized variants, the learnable radial function $R^{(t)}_{k\eta_{1}l_{1}l_{2}l_{3}}(r_{ij})$ is evaluated with cubic splines, which efficiently map $r_{ij}$ to each of the outputs of the radial MLP. Furthermore, eqs. II.1 and 3 are implemented in unified kernels that simultaneously compute the one-particle basis and perform the summation over neighbors to generate the invariant atomic basis $A_{i}$. The computation of spherical harmonics, as detailed in equation II.1 is accelerated using the sphericart [90] library, which offers a highly optimized implementation for real spherical harmonics. All linear and equivariant linear operations have been substituted with custom kernels, and the cuda_mace implementation further leverages tensor core acceleration while incorporating error-correction techniques to ensure numerical accuracy. Additionally, the computation of the product basis $\mathbf{B}_{i}$ in eq 4 is optimized using the sparsity of the generalized Clebsch-Gordon coefficients $\mathcal{C}^{LM}_{\eta_{\nu}\bm{l}\bm{m}}$, significantly accelerating the product calculations over both the body order $v$ and the coupling terms.
@@
Figure 12: Performance in OpenMM. Performance of MACE-OFF models in OpenMM for liquid water in the NVT ensemble at 1 g/cm3 at 300 K with a 1 fs timestep on a single NVIDIA 80GB A100 GPU. The cuda_mace implementation is labeled ‘OPT’. Forces were evaluated in double precision, and integrated in single precision.
@@ IV Outlook
In this paper, we presented the series of MACE-OFF force field models, demonstrating the broad applicability, transferability and capability of purely short-range machine learning potentials for organic (bio)molecular simulations. We have shown that the models, based on the MACE higher order equivariant message passing architecture, can improve on the pioneering ANI models [27] that served as the only widely applicable machine learning force fields for molecules for a number of years. While a full ablation study between different ML models is not possible due to a wide range of design choices made in their construction, the MACE-OFF model architecture, training data and loss function all contribute to significant improvements in both accuracy and extrapolation, combined with high computational speed.
We trained here so-called small, medium and large versions of the MACE-OFF models, with systematically improving accuracy when benchmarked against gas phase quantum mechanical data. We further investigated the accuracy of the medium models against a range of crystalline, liquid and biomolecular condensed phase data, as we found them to provide a good compromise between accuracy and computational expense. We have made all force fields available to the community, so that users can choose the model most suited to their application.
@@ Acknowledgements
The data used to train the models are publicly available at: https://doi.org/10.17863/CAM.107498. The torsion drive dataset is also available at: https://zenodo.org/records/11385284. The MACE-OFF series of models is available at: https://github.com/ACEsuit/mace-off/tree/main.